A 40g body falls, from rest, through a vertical distance of 25 cm to the ground. What is the kinetic energy and the velocity of the body just before hitting the ground?
Answer #1

A 40g body falls, from rest, through a vertical distance of 25 cm to the ground. What is the kinetic energy and the velocity of the body just before hitting the ground?


Distance, s = 25 cm = 0.25 meter
Mass of body, m = 40 g = 40/1000 kg = 0.04 kg
Initial velocity, u = 0 m/s (dropped from rest)
acceleration, a = 9.8 m/s2

Potential energy is converted to kinetic energy,

mgh = 1/2mv^2
sqrt is square root
v = sqrt(2gh) = sqrt(2*9.8*0.25) = 2.2136 m/s
Kinetic energy = 0.04*9.8*0.25 = 0.098 Joule
Speed = 2.2136 m/s